A body is projected horizontally with 10 m/s from the top of a
tower of height 20 m, on reaching the ground its displacement is
(g = 10 m/s^2)
D)20 m
E)40 m
F)40 √ 2 m
G)20 √ 2 m


 

Grade:11

1 Answers

Vikas TU
7 years ago
Apply the eqns. in horizontally and vertical respectively.
In x – direcn. we get,
x = 10*t
In y- direcn. we get,
 
20 = 0 + 0.5gt^2
t  = 2 seconds.
Hence,
x = 10*2 = 20 meter.